KjSaif (kjsaif):

Solve for x. 1/2 x =20 140 110 10 40

OpenStudy (divyankasc):

1/2 × x = 10 That means x/2 = 10 We now take the denominator 2 on LHS to RHS What we get is this → x = 2 × 10 x = 40!

OpenStudy (divyankasc):

x = 40*

KjSaif (kjsaif):

thanks

OpenStudy (divyankasc):

Welcome :)